Big plane. A six-passenger plane cruises at 180 mph in calm air. If the plane flies 7 miles with the wind in the same amount of time as it flies 5 miles against the wind, then what is the wind speed?

Answer by checkley77(12844)

D-RT
7=(280+W)T
T=7/(180+W)FORMULA FOR THE TRIP WITH THE WIND.
5=(180-W)T
T=5(180-W) FORMULA FOR THE TRIP AGAINST THE WIND.
SET THESE 2 T FORMULAS EQUAL
7/(180+W)=5/(180-W) CROSS MULTIPLY.
5(180+W)=7(180-W)
900+5W=1,260-7W
5W+7W=1,260-900
12W=360
W=360/12
W=30 MPH. IS THE WUIND SPEED.
PROOF:
7/(180+30)=5/(180-30)
7/210=5/150
1/30-1/30